# Pulsed magnetic field gradient on a tip for nanoscale imaging of spins

### v1

@@ Methods
The process for fabricating the current carrying tip is described in the following. The base material is a 100 mm-long quartz solid rod (no inner diameter) with two grooves along opposite sides (see Extended Figure 1). The rod is pulled to form two tips using Sutter Instruments’ P-2000, which heats the rod at the center and pulls the two tapers to form two tips. The tips have an apex on the order of 100 nm and taper length on the order of 5 mm, where the precise dimensions depend on the working parameters. A home-built thermal deposition [31] is used to deposit 2 nm Cr followed by 10 nm Au on the two long sides of the rod, creating two leads. To form a 1 apex, a focused (gallium) ion beam was used to cut the apex by removing the quartz. A second metal deposition was then performed on the two leads and the apex: At the apex, 7 nm Cr and 120 nm Au were deposited and on each of the leads 7 nm Cr and 50 nm Au.
@@
The tip is placed on a piezo stage which controls its location in x, y and z relative to the diamond sample. An elliptical tip holder was designed to hole the tip at the center with two metal springs that provide electric contact to each lead. Each spring is connected to a coaxial cable to connect to the electric current circuit. A quartz tuning fork (, HM International BVBA, model TB-38-20-12.5-32.768) is pressed against the tip using a micropositioner and a live camera, ensuring tip to tuning fork contact and preventing the tuning fork from damaging the tip.
@@
Supplementary information Effect of increasing current on the measured gradient, assessing the gradient in Rabi frequency, simulating the device heating, modelling the effect of a metallic tip on the Rabi driving power.
Acknowledgements We are thankful to D. Yudilevich for providing us with an in-depth proof of the manuscript. We thank E. Zeldov and Y. Myasoedov for giving us access to their tip puller and thermal deposition tool. We are grateful to U. Goldblatt, F. Lafont and S. Rosenblum for assistance with high-frequency microwave simulations (HFSS). A. F. is the incumbent of the Elaine Blond Career Development Chair in Perpetuity. We acknowledge support by the Kimmel Institute for Magnetic Resonance. This research is made possible in part by the historic generosity of the Harold Perlman Family,
Declarations
Funding
This research was funded exclusively from internal Weizmann Institute of Science resources. -
@@
The authors declare no competing interests. The data relevant to figures in the main text are available via the Weizmann/Elsevier data repository, DOI:10.34933/TBD. Additional raw data are available from the corresponding author upon reasonable request. -
Code availability The MATLAB code is available from the authors upon reasonable request.

### v2

@@ Methods
The process for fabricating the current carrying tip is described in the following. The base material is a 100 mm-long quartz solid rod (no inner diameter) with two grooves along opposite sides (see Supplementary Note 2). The rod is pulled to form two tips using Sutter Instruments’ P-2000, which heats the rod at the center and pulls the two tapers to form two tips. The tips have an apex on the order of 100 nm and taper length on the order of 5 mm, where the precise dimensions depend on the working parameters. A home-built thermal deposition [38] is used to deposit 2 nm Cr followed by 10 nm Au on the two long sides of the rod, creating two leads. To form a 1 apex, a focused (gallium) ion beam was used to cut the apex by removing the quartz. A second metal deposition was then performed on the two leads and the apex: At the apex, 7 nm Cr and 120 nm Au were deposited and on each of the leads 7 nm Cr and 50 nm Au.
@@
The tip is placed on a piezo stage which controls its location in x, y and z relative to the diamond sample. An elliptical tip holder was designed to hold the tip at the center with two metal springs that provide electric contact to each lead. Each spring is connected to a coaxial cable to connect to the electric current circuit. A quartz tuning fork (, HM International BVBA, model TB-38-20-12.5-32.768) is pressed against the tip using a micropositioner and a live camera, ensuring tip to tuning fork contact and preventing the tuning fork from damaging the tip.
@@
Data Availability
The data relevant to figures in the main text are available via the Weizmann/Elsevier data repository, DOI: 10.34933/f3595abc-d114-4ee3-9be1-0739d7d1360d [39]. Additional raw data are available from the corresponding author upon reasonable request.
Code Availability
The MATLAB code is available from the authors upon reasonable request from the corresponding author.
Acknowledgments
We are thankful to D. Yudilevich for providing us with an in-depth proof of the manuscript. We thank E. Zeldov and Y. Myasoedov for giving us access to their tip puller and thermal deposition tool. We are grateful to U. Goldblatt, F. Lafont and S. Rosenblum for assistance with high-frequency microwave simulations (HFSS). A. F. is the incumbent of the Elaine Blond Career Development Chair in Perpetuity. We acknowledge support by the Kimmel Institute for Magnetic Resonance. This research is made possible in part by the historic generosity of the Harold Perlman Family. This research was funded exclusively from internal Weizmann Institute of Science resources.
@@ Acknowledgments
The authors declare no competing interests.